OpenStudy (anonymous):

lim n→∞ (sqrt{n2−1}−n.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\sqrt{n²-1} -n\]

OpenStudy (experimentx):

multiply the expression by it's conjugate

OpenStudy (anonymous):

\[\frac{ (\sqrt{n²−1}−n)(\sqrt{n²−1}+n) }{ (\sqrt{n²−1}+n)}\]

OpenStudy (experimentx):

yeah .. that's going on right direction.

OpenStudy (anonymous):

yeah sry writing eqations with the platfom is so dam slow.. I have the solution, but I don't understand it. first question would be why the numerator =-1 ?

OpenStudy (experimentx):

doesn't numerator look like (a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

aha. you're right

OpenStudy (anonymous):

thanks @experimentX

OpenStudy (experimentx):

yw